A motorist travelling with a constant speed 15m/s passes a school crossing corner where speed limit is 10m/s just as the motorist passes a police on a motorcycle starts off in pursuit with a constant acceleration 3m/s2 .How much time elapses before he catches up with the motorist?

What is the officers speed? What is the total distance travelled byveach vehicle at that point?


1
Expert's answer
2021-03-21T11:20:55-0400

(a) The distance traveled by the motorist can be found as follows:


"d_1=v_0t=15t."

The distance traveled by the police on a motorcycle can be found as follows:


"d_2=\dfrac{1}{2}at^2=1.5t^2."

When the police on the motorcycle catches up with the motorist, "d_1=d_2":


"15t=1.5t^2,""t=\dfrac{15\ \dfrac{m}{s}}{1.5\ \dfrac{m}{s^2}}=10\ s."

(b) We can find the officer's speed as follows:


"v=v_0+at=0+3\ \dfrac{m}{s^2}\cdot10\ s=30\ \dfrac{m}{s}."

(c) We can find the total distance travelled by each vehicle at that point as follows:


"d=v_0t=15\ \dfrac{m}{s}\cdot10\ s=150\ m."